OpenStudy (anonymous):

How does changing the function from f(x) = 3 sin 2x to g(x) = 3 sin 2x + 5 affect the range of the function? A. The function shifts up 3 units, so the range changes from −1 to 1 in f(x) to 2 to 4 in g(x). B. The function shifts up 3 units, so the range changes from −3 to 3 in f(x) to 0 to 6 in g(x). C. The function shifts up 5 units, so the range changes from −1 to 1 in f(x) to 4 to 6 in g(x). D. The function shifts up 5 units, so the range changes from −3 to 3 in f(x) to 2 to 8 in g(x).

OpenStudy (anonymous):

@oleg3321 @kropot72

OpenStudy (anonymous):

@xapproachesinfinity

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Ibbutibbu.

OpenStudy (ibbutibbu.):

yeah?

OpenStudy (anonymous):

can you please help me

OpenStudy (ibbutibbu.):

i thinks itd D

OpenStudy (anonymous):

thanks

OpenStudy (ibbutibbu.):

np

OpenStudy (xapproachesinfinity):

ok, let me explain

OpenStudy (xapproachesinfinity):

g(x)=f(x+L) L units to left g(x)=f(x-L) L units to right g(x)=f(x)+L L units UP g(x)=f(x)-L L units down

OpenStudy (xapproachesinfinity):

can you see one of these in this problem?

OpenStudy (anonymous):

thanks @xapproachesinfinity I understand know, do u mind helping me out with another one?

OpenStudy (xapproachesinfinity):

go ahead